OpenStudy (anonymous):

Jonah was keeping track of how many cars drove by every hour. He realized that the number of cars were in a sequence that followed the pattern n +3. How many cars drove by on the 9th hour? 27 6 12 3

OpenStudy (hoblos):

U(n) = n+3 so U(9) = 9+3 = 12